How would you set up the following word problem:
You are hitting baseballs. When tossing the ball into the air, your hand is 5 feet above the ground. You hit the ball when it falls back to a height of 4 ft. If you toss the ball with an initial velocity of 25ft/sec, how much time will pass before you hit the ball?
Thanks!

Answer by AnlytcPhil(1810)

How would you set up the following word problem: You are
hitting baseballs. When tossing the ball into the air, your
hand is 5 feet above the ground. You hit the ball when it
falls back to a height of 4 ft. If you toss the ball with
an initial velocity of 25ft/sec, how much time will pass
before you hit the ball? Thanks!
Use the formula h = h0 + v0t - 16t2 since distance is given
in feet. (If distance were in meters, we would have to use
4.9 where the 16 is).
h0 represents the initial height. v0 represents the initial
velocity, and h is height at the instant when t seconds have
elapsed.
>>...When tossing the ball into the air, your hand is 5 feet
above the ground...<<
So h0 = 5
>>...you toss the ball with an initial velocity of 25ft/sec...<<
So v0 = 25
So we substitute 5 for h0 and 25 for v0, making our formula
h = 5 + 25t - 16t2
>>...You hit the ball when it falls back to a height of 4 ft...<<
So we wish to find t (the number of seconds) has elapsed at the
instant when the ball is 4 feet high, so we substitute 4 for h
and solve for t:
4 = 5 + 25t - 16t2
16t2 - 25t - 1 = 0
Use quadratic formula, get solutions
-.04 and 1.6. We ignore the negative answer. So the answer is
1.6 seconds approximately.
Edwin